40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Oct 29 13:48:32 2009, Koji, Update, PSL, NPRO LTMP lowered 9.5deg 091028_PSL.png
    Reply  Thu Oct 29 20:21:14 2009, Koji, Update, PSL, NPRO LTMP lowered 9.5deg PSL_MC.png
       Reply  Fri Oct 30 09:24:45 2009, steve, HowTo, MOPA, how to squeeze more out of little 44to34.jpg
Message ID: 2158     Entry time: Thu Oct 29 13:48:32 2009     Reply to this: 2161
Author: Koji 
Type: Update 
Category: PSL 
Subject: NPRO LTMP lowered 9.5deg 

13:00 Found MC TRANS less than 7.
13:50 Go into the PSL table.
14:20 Work done. Now I am running SLOWscan script.
15:10 SLOWscan finished. It was not satisfactory. I go into the table again.
15:15 Running SLOWscan again.
16:00 SLOWscan done. Lock PMC. Adjust NPRO current so as to maximize PMC TRANS.
16:10 Lock RC, PMC, MZ, MC. Align PMC / MZ on the table. Align MC WFS beams on the QPDs.
16:30 Work done.

New FSS-SLOWDC nominal is -4.0

Now MC TRANS is 7.9. This is +12% increase. ENJOY!
HEPA is on at 90%. Light is off.

---------

NPRO TEMP trimmer adjustment
o PSL NPRO TEMP trimmer at the back of the laser head was turned 6.5 times in CW.
o It reduced NPRO crystal temp by 9.5deg. (43.5deg -> 34.0deg for FSS_SLOWDC -5.5)

To revert the previous setting, refer to the former measurement
c.f. http://nodus.ligo.caltech.edu:8080/40m/2008

NPRO Thermal scan
o 2 scans are performed.
o I selected the colder side of the second scan. i.e. SLOWDC=-4.0

NPRO Current adjustment
o Tweaked C1:PSL-126MOPA_126CURADJ while looking at PMC TRANS.
o CURADJ was changed from -2.25 to -1.9. This corresponds to change of C1:PSL-126MOPA_CURMON from 2.503A to 2.547A.

Attachment 1: 091028_PSL.png  29 kB  Uploaded Thu Oct 29 19:02:27 2009  | Hide | Hide all
091028_PSL.png
ELOG V3.1.3-